OpenStudy (anonymous):

lim as x approaches 0 of (x^2)*(3+sin(x))/((x+sin(x))^2) I know I'm supposed to use the fact that lim as x approaches 0 of sin(x)/x = 1

OpenStudy (jamesj):

Divide top and bottom by x^2. Then I think you'll see it clearly.

OpenStudy (anonymous):

ok I'll try that and see what happens

OpenStudy (jamesj):

got it?

OpenStudy (anonymous):

I've got (3+sin(x))/(x+sin(x)/x)^2 but I still can't see how to get the sin(x)/x out of it

OpenStudy (jamesj):

your denominator isn't correct. check it again carefully

OpenStudy (anonymous):

isn't (x+sin(x)/x)^2 equal to (x+sin(x))^2/x^2

OpenStudy (jamesj):

no ...

OpenStudy (anonymous):

should I expand the (x+sin(x))^2?

OpenStudy (jamesj):

\[\frac{1}{x^2} (x + \sin x)^2 = \left( x/x + \sin x / x \right)^2 = (1 + \sin x/x)^2\]

OpenStudy (anonymous):

Oh! I distributed the \[1/x^2\] wrong

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (jamesj):

so what's the answer?

OpenStudy (anonymous):

3/4 I think

OpenStudy (jamesj):

Yes, that's right.

OpenStudy (anonymous):

Thanks for helping me :)